# Towards Visual Text Design Transfer Across Languages

### v1

@@ Abstract
The art of visual text design serves as a potent medium for conveying themes, emotions, and atmospheres within a multimodal context. From compelling film posters to evocative album covers, the fusion of typography and imagery transcends the communicative potential of mere words. Nevertheless, the translation of a visual style’s essence across disparate writing systems presents a substantial challenge for computational models. Can generative models accurately comprehend the intricacies of design and effectively transfer the intended aesthetic across linguistic boundaries? In this study, we introduce Multimodal Style Translation (MuST-Bench), a pioneering task designed to evaluate the efficacy of visual text translation across diverse writing systems. Our studies with MuST-Bench reveal that current visual text generation models struggle with the proposed task due to the inadequacy of textual descriptions in conveying visual design. We introduce SIGIL, a framework for multimodal style translation that eliminates the need for style descriptions. SIGIL enhances image generation models through three innovations: glyph latent for multilingual settings, pretrained VAEs for stable style guidance, and an OCR model with reinforcement learning feedback for optimizing readable character generation. SIGIL surpasses baselines in style consistency and legibility while maintaining visual similarity, unlike description-based methods. We plan to release our benchmark and model to inspire further research in multilingual visual text understanding and generation.
11footnotetext: denotes equal contributionfootnotetext: Data: https://huggingface.co/datasets/yejinc/MuST-Bench
@@ 1 Introduction
Recent advancements in image generation have greatly improved the creation of visual text, allowing for instruction-based control over text style. However, these text-to-image models often fall short of accurately conveying abstract concepts like "design" and "style" through text alone. In practical scenarios, hand-crafted artistic visual text typically lacks precise descriptions of these concepts. For instance, in film posters, album covers, or advertisements, human experts are essential to communicate the main theme and atmosphere. This raises an important question: can we evaluate a machine’s ability to understand and transfer such diverse concepts in a multimodal context? To address this, we introduce a novel task called multimodal style translation, specifically designed for controlled evaluation of creative visual text style generation. We define multimodal style translation as a description-free, few-shot style transfer task across different writing systems, challenging models to capture the essence of style design beyond just the font’s surface form.
MuST-Bench, our proposed benchmark, is a carefully curated dataset that extends beyond simple multilingual data collection by comparing language pairs with similar stylistic features across diverse typographical styles. This ensures a robust evaluation of models’ capabilities in capturing the nuances of visual text design across different linguistic and cultural contexts. Furthermore, each dataset entry is augmented with human-annotated bounding boxes at the character level, substantially enhancing the benchmark’s quality and precision. MuST-Bench samples feature natural artistic typographies sourced from film posters, enabling the evaluation of transfer capabilities from English to five target languages with diverse writing systems: Chinese, Korean, Thai, Russian, and Arabic. However, to the best of our knowledge, no previous methods have been proposed to handle multilingual visual text generation conditioned on a few-shot image samples. Additionally, as we empirically demonstrate later (see section 5.2), state-of-the-art visual text generation models struggle with our benchmark, failing to capture the intricate designs present in artistic typography instances.
To address the challenges, we introduce the Style Integrity and Glyph Incentive Learning ([Uncaptioned image]SIGIL ) framework. SIGIL introduces two technical novelties: first, SIGIL guides the character generation process using a distance metric defined in the pretrained glyph latent space. This space, based on a pretrained VAE, allows for the comparison of glyphs across different languages. Our empirical results demonstrate that guidance within the glyph latent space achieves superior style fidelity compared to textual conditioning. Second, we employ an OCR model to generate confidence scores for the images during the generation process. These scores serve as rewards in a reinforcement-learning-based optimization, enhancing the readability of the generated text.
@@ 2 Related Work
Recent advancements in text-to-image generation tasks [3, 7, 19], notably with models like DALLE-3 [3], have demonstrated exceptional capabilities in image synthesis. Despite these advancements, generating precise and coherent text within images remains challenging. To improve text coherence and accuracy, models like Imagen [19] and DeepFloyd IF [6] employ pretrained large language models such as T5-XXL [16]. Stable Diffusion 3  [7] introduces a novel transformer-based architecture that uses separate weights for image and text tokens, enhancing text comprehension. However, these methods require an extensive image-caption pair dataset, making it difficult to extend them to languages other than English.
@@ 4.2 Glyph Guidance on the Latent Space
We adopt a latent diffusion setting [17], where the diffusion process occurs in the latent space $z$ defined by a pretrained variational autoencoder (VAE) [10] encoder, rather than in the pixel space.
@@
By using the latent semantic space as our metric space, we provide a cost-effective way to compare characters across languages without training a separate classifier. As previous literature suggests [27, 12], representation spaces defined by pretrained neural networks often better correlate with human perception. By using the same VAE for both diffusion framework and guidance loss, we provide an efficient solution for controlling glyph shapes. We demonstrate the effectiveness of latent space distances through qualitative ablation studies in fig. 7.
@@ 5.2 Quantitative Evaluation
For the style fidelity evaluation using MLLMs, we employed the method mentioned in section 3. To ensure the reproducibility of the evaluation results, we used a consistent input judgment template. For GPT-4V, we set the seed to 100 and the temperature to 0. Although Claude-3-OPUS does not support seed fixation, setting the temperature to 0 minimizes variability. To further ensure reproducibility, we provided consistent input data and verified the results across multiple runs. For each evaluation, the images generated by each method were randomly shuffled before being presented to the MLLM for assessment. A total of 30 questions were involved in the evaluation process. The evaluation results are presented in table 3. SIGIL achieved the highest style fidelity scores on both GPT-4V and Claude. Interestingly, even though images were generated from DALL-E3 using prompts created by GPT-4V based on the input style image, SIGIL still received higher scores in the GPT-4V evaluation. During the evaluation, we identified a safety misclassification issue with GPT-4V for some typographic images, and four such samples are included in Appendix A.

### v2

@@ Abstract
Visual text design plays a critical role in conveying themes, emotions, and atmospheres in multimodal formats such as film posters and album covers. Translating these visual and textual elements across languages extends the concept of translation beyond mere text, requiring the adaptation of aesthetic and stylistic features. To address this, we introduce a novel task of Multimodal Style Translation (MuST-Bench), a benchmark designed to evaluate the ability of visual text generation models to perform translation across different writing systems while preserving design intent. Our initial experiments on MuST-Bench reveal that existing visual text generation models struggle with the proposed task due to the inadequacy of textual descriptions in conveying visual design. In response, we introduce SIGIL, a framework for multimodal style translation that eliminates the need for style descriptions. SIGIL enhances image generation models through three innovations: glyph latent for multilingual settings, pre-trained VAEs for stable style guidance, and an OCR model with reinforcement learning feedback for optimizing readable character generation. SIGIL outperforms existing baselines by achieving superior style consistency and legibility while maintaining visual fidelity, setting itself apart from traditional description-based approaches. We release MuST-Bench publicly for broader use and exploration [^1] https://huggingface.co/datasets/yejinc/MuST-Bench.
11footnotetext: denotes equal contribution
@@ 1 Introduction
Recent advancements in image generation have significantly enhanced the ability to create and manipulate text within images, enabling instruction-based control over text styles. However, achieving this level of fine-grained style manipulation becomes challenging in cross-language contexts. For instance, imagine the title text on a sci-fi movie poster, initially rendered in English with a futuristic, metallic font, being translated to Korean while preserving the same font style, color gradients for visual impact. Effective multimodal translation requires two factors: transferable visual content across languages and accurate multilingual text generation that maintains the original design.
To tackle this issue, we introduce MuST-Bench, a carefully curated dataset that compares language pairs with similar stylistic features across diverse typographical styles. This dataset enables robust evaluation of models’ abilities to capture visual text design nuances across languages. Furthermore, each dataset entry is augmented with human-annotated bounding boxes at the character level, substantially enhancing the benchmark’s quality and precision. MuST-Bench samples feature natural artistic typographies sourced from film posters, enabling the evaluation of transfer capabilities from English to five target languages with diverse writing systems: Chinese, Korean, Thai, Russian, and Arabic.
However, current text-to-image models generally rely on textual descriptions for image generation, which restricts their effectiveness in practical visual contexts such as film posters, album covers, or advertisements, where meticulously crafted artistic composition is crucial to communicate the main theme and atmosphere. In such cases, abstract concepts as design and style are hard-explained through text alone. As we empirically demonstrate later (see section 5.2), state-of-the-art visual text generation models struggle with our benchmark, failing to capture the intricate designs present in artistic typography instances.
To address the challenges, we introduce the Style Integrity and Glyph Incentive Learning ([Uncaptioned image]SIGIL ) framework. SIGIL introduces two technical novelties: First, SIGIL guides the character generation process using a distance metric defined in the pre-trained glyph latent space. This space, based on a pre-trained VAE, allows for the comparison of glyphs across different languages. Our empirical results demonstrate that guidance within the glyph latent space achieves superior style fidelity compared to textual conditioning. Second, we employ an OCR model to generate confidence scores for the images during the generation process. These scores serve as rewards in a reinforcement-learning-based optimization, enhancing the readability of the generated text.
@@ 2 Related Work
Recent advancements in text-to-image generation tasks [3, 7, 19], notably with models like DALLE-3 [3], have demonstrated exceptional capabilities in image synthesis. Despite these advancements, generating precise and coherent text within images remains challenging. To improve text coherence and accuracy, models like Imagen [19] and DeepFloyd IF [6] employ pre-trained large language models such as T5-XXL [16]. Stable Diffusion 3  [7] introduces a novel transformer-based architecture that uses separate weights for image and text tokens, enhancing text comprehension. However, these methods require an extensive image-caption pair dataset, making it difficult to extend them to languages other than English.
@@ 4.2 Glyph Guidance on the Latent Space
We adopt a latent diffusion setting [17], where the diffusion process occurs in the latent space $z$ defined by a pre-trained variational autoencoder (VAE) [10] encoder, rather than in the pixel space.
@@
By using the latent semantic space as our metric space, we provide a cost-effective way to compare characters across languages without training a separate classifier. As previous literature suggests [27, 12], representation spaces defined by pre-trained neural networks often better correlate with human perception. By using the same VAE for both diffusion framework and guidance loss, we provide an efficient solution for controlling glyph shapes. We demonstrate the effectiveness of latent space distances through qualitative ablation studies in fig. 7.
@@ 5.2 Quantitative Evaluation
For the style fidelity evaluation using MLLMs, we employed the method mentioned in section 3. To ensure the reproducibility of the evaluation results, we used a consistent input judgment template. For GPT-4V, we set the seed to 100 and the temperature to 0. Although Claude-3-OPUS does not support seed fixation, setting the temperature to 0 minimizes variability. To further ensure reproducibility, we provided consistent input data and verified the results across multiple runs.
For each evaluation, the images generated by each method were randomly shuffled before being presented to the MLLM for assessment. A total of 30 questions were involved in the evaluation process. The evaluation results are presented in table 3. SIGIL achieved the highest style fidelity scores on both GPT-4V and Claude. Interestingly, even though images were generated from DALL-E3 using prompts created by GPT-4V based on the input style image, SIGIL still received higher scores in the GPT-4V evaluation. During the evaluation, we identified a safety misclassification issue with GPT-4V for some typographic images, and four such samples are included in Appendix A.